OpenStudy (anonymous):

Assume that y varies inversely as x. If y = 10 when x = 5, find y when x = 2.

OpenStudy (jhannybean):

inverse variation follows the format y= k/x you're given y=10 ,x=5 so first we find value of k. 10 = k/ (5) k = 5(10) = 50. Now we're trying to find y when x=2 so you use the same formula,but this time use k=50 and x=2. y = k/x y = 50 / 2 y = 25